OpenStudy (anonymous):

David has a bag that holds 18 pens. 5 of the pens are ball point, 4 are gel pens, and 9 are felt tip marker pens. He reaches into the bag to pull out one pen. If every pen has an equal chance of being pulled out of David’s bag, what is the probability that he will pull out a gel pen

OpenStudy (anonymous):

it is the ratio of the number of gel pens to the total number of pens

OpenStudy (anonymous):

how do i figure it out?

OpenStudy (anonymous):

how many gel pens are there?

OpenStudy (anonymous):

altogether?

OpenStudy (anonymous):

yes how many gel pens all together in the bag?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

actually that is the total number of pens of those, how many are gel pens?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

right your answer is \[\frac{4}{18}=\frac{2}{9}\]

OpenStudy (anonymous):

i.e. \[\frac{\text{number of gel pens}}{\text{total number of pens}}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw say hi to randy for me

OpenStudy (anonymous):

haha okayy!!!! :)